# Rediscovering Entropy Regularization: Adaptive Coefficient Unlocks Its Potential for LLM Reinforcement Learning

### v1

@@ C.1 Pass@$k$ as a Function of $k$ (Qwen3 - 8B - Base)
AER ($\beta{=}0.2$) AIME2024 AIME2025 AMC23 MATH500 Avg.
@@
We ablate the difficulty threshold $\rho$ (formerly denoted $\beta$; cf. Section §4.2) while fixing the target-entropy ratio at $\tau{=}0.4$. Recall that samples with group accuracy $g(q)\leq\rho$ receive a positive entropy bonus (hard set). Table 8 shows that narrower hard sets (smaller $\rho$) generally yield higher overall accuracy: the best average is obtained at $\rho{=}0.0$ (55.8), and performance degrades when $\rho$ is made overly inclusive (e.g., $\rho{=}1.0$, 53.5). This indicates that allocating the entropy budget only to the genuinely difficult prompts preserves exploitation on easier ones and avoids diluting exploration across too many samples. At the dataset level, AIME2025 and AMC23 achieve their peaks at moderate $\rho$ (0.4/0.8 and 0.4, respectively), suggesting a mild benefit from targeted—but not indiscriminate—expansion of the hard set.

### v2

@@ C.1 Pass@$k$ as a Function of $k$ (Qwen3 - 8B - Base)
AER ($\rho{=}0.2$) AIME2024 AIME2025 AMC23 MATH500 Avg.
@@
We ablate the difficulty threshold $\rho$ (formerly denoted $\rho$; cf. Section §4.2) while fixing the target-entropy ratio at $\tau{=}0.4$. Recall that samples with group accuracy $g(q)\leq\rho$ receive a positive entropy bonus (hard set). Table 8 shows that narrower hard sets (smaller $\rho$) generally yield higher overall accuracy: the best average is obtained at $\rho{=}0.0$ (55.8), and performance degrades when $\rho$ is made overly inclusive (e.g., $\rho{=}1.0$, 53.5). This indicates that allocating the entropy budget only to the genuinely difficult prompts preserves exploitation on easier ones and avoids diluting exploration across too many samples. At the dataset level, AIME2025 and AMC23 achieve their peaks at moderate $\rho$ (0.4/0.8 and 0.4, respectively), suggesting a mild benefit from targeted—but not indiscriminate—expansion of the hard set.